OpenStudy (ujjwal):

Find the equation of plane perpendicular to the yz plane and passing through points (2,3,1) and (4,-5,3)

OpenStudy (experimentx):

let the plane be ... ax+by+cz+d=0 put the values of x and y, 2a+3b+c+d=0 4a-5b+3c+d=0 and for perpendicular planes, normals are perpendicular <a,b,c><0,0,1>=0 c=0

OpenStudy (anonymous):

oh i dont think i is n.v ,sry...let me see..

OpenStudy (experimentx):

i think, more simpler .... 2a+3b+c=1 4a-5b+3c=1

OpenStudy (anonymous):

passing through points (2,3,1) and (4,-5,3) means the vector v(2,-8,2) is paralel to the plane.Equation of yz plane is : x=0. So perpendicular vector to it is:p(1,0,0) B=vXp this is the vector perpendicular to the plane you looking for (x represents vector product)

OpenStudy (experimentx):

and normal of yz plane  I think it should be like this